# Active Inference Meeting Energy-Efficient Control of Parallel and Identical Machines

### v1

@@ Abstract
Keywords Active Inference  $\cdot$ Probabilistic Deep Learning  $\cdot$ Reinforcement Learning  $\cdot$ Energy-Efficient Control  $\cdot$ Manufacturing Systems
@@ 1 Introduction
Active inference (AIF), an emerging field inspired by the principles of biological brains, offers a promising alternative for decision-making models. It unifies perception, learning, and decision-making under the free energy principle (FEP), which formulates neuronal inference and learning under uncertainty [1]. Accordingly, the brain is modeled through levels of (variational) Bayesian inference [2], minimizing prediction errors by leveraging a generative model of the world while considering uncertainties. This framework enables the development of agents that can calibrate their models and make decisions without complete knowledge of system dynamics. Significant progress has been made in applying active inference across various domains, including robotics, autonomous driving, and healthcare [3, 4, 5], showcasing its ability to handle complex decision-making tasks in dynamic environments.
Recently, the manufacturing industry’s focus on energy efficiency has intensified due to its significant contribution to global energy consumption. Addressing energy efficiency at the machine level has become critical, with energy-efficient scheduling (EES) and energy-efficient control (EEC) strategies emerging as key approaches to reducing environmental impact [6]. While traditional EEC methods often necessitate complete system knowledge, reinforcement learning (RL) has shown potential in optimizing manufacturing processes without prior system knowledge [7, 8]. However, RL agents may struggle to rapidly adjust their policies to changing conditions.
@@ 2 Application Overview
2.1 EEC in Manufacturing Systems
EEC gaining prominence in both academic and industrial circles within manufacturing systems. It offers substantial energy savings across three key control levels: component, machine, and production system. At its core, EEC involves managing the power consumption state of objects based on environmental conditions. Objects are kept fully operational when their functions are needed and transitioned to low power states when not in use, though this poses challenges due to the unpredictable nature of environmental demands and the penalties incurred during state transitions. These penalties include both time wasted during transition, when the object adds no value, and the energy consumed during the transition process. A comprehensive and recent literature review on this topic can be found in [15].
2.2 System Description
The specific system under study is a manufacturing workstation consisting of a single upstream buffer serving multiple parallel machines as depicted in Fig. 1. Key characteristics of this system include a finite buffer capacity, stochastic arrival of parts, and identical machines operating in a mass-production context. Each machine can exist in various states including working, standby, startup, and failed, with further subdivisions within the working state to denote idle and busy states. Additionally, machines are subject to stochastic failures and repairs, each with expected durations, further complicating control strategies.
Figure 1: Layout of parallel machines.
Power consumption varies based on the machine’s state, with different power requirements for standby, startup, idle, and busy states. For instance, standby and failed states require minimal power consumption, while idle and startup states demand higher levels. Furthermore, the highest power consumption occurs during the busy state when the machine is fully operational and actively processing parts. Managing these power variations efficiently is crucial for effective energy-efficient control in manufacturing systems.
@@ 3.1 Active Inference Formalism
where $\sigma(\cdot)$ represents the Softmax function, and $\pi$ (i.e., policy) denotes the sequence of actions. The EFE encompasses minimizing surprise regarding preferred observations [^1] Here, the surprise refers to the preference, while in VFE, surprise pertains to the actual observation used for calibrating the model., exploring uncertainty, and reducing uncertainty about model parameters [9]. The EFE for $\tau\geq t$ can be formulated as follows [25]:
@@ 3.2 Architecture
An agent within the active inference framework needs different modules, which are entangled within the framework. Amortization [26, 27] is introduced into the formalism (in Sec. 3.1) to scale-up the realization [9]. The formalism is inherently probabilistic and parameterized with with two sets, $\theta=\{{\theta}_{s},{\theta}_{o}\}$ representing generative and $\phi=\{{\phi}_{s}\}$ recognition elements [9]. Using the following parameterized modules the formalism (in Sec. 3.1) can be calculated: Encoder (i.e., $Q_{\phi_{s}}(s_{t})$), an amortized inference of the hidden state (i.e., an inference network [28] providing a mapping between the observation, $\tilde{o}_{t}$, and a distribution for its corresponding hidden state). Transition (i.e., ${P}_{\theta_{s}}({s}_{t+1}|\tilde{s}_{t},\tilde{a}_{t})$), which generates a distribution for the next hidden state based on both a sampled action and the current hidden state. Decoder (i.e., ${P}_{\theta_{o}}({o}_{t+1}|\tilde{s}_{t+1})$) generates a distribution for prediction based on the sampled hidden state.
@@
Utilizing the mentioned architecture (also depicted in Fig. 3), the agent would be able to calculate the EFE in Eq. 5 for a given policy (i.e., $\pi$), which is a sequence of actions. Thus, we can calculate the probabilities for selecting actions through Eq. 2. In order to make better decisions, the agent can plan ahead by simulating future trajectories using the architecture mentioned above. However, as the policy space grows exponentially into the future, it is infeasible to evaluate all possible scenarios. Fountas et al. (2020) [9] introduced two approaches to alleviate the obstacle. 1) They used the standard Monte-Carlo Tree Search (MCTS) [29, 30], a tree-based search algorithm, which selectively explore promising trajectories in a restricted manner. 2) They introduced another recognition module [31, 32, 27], parameterized with ${\phi}_{a}$ as follows: Habit (i.e., $Q_{\phi_{a}}(a_{t})$), an amortized inference of actions (i.e., an inference network [28] providing a mapping between a sampled hidden state, $\tilde{s}_{t}$, and normalized, through Softmax function, probabilities for the actions). Habit is also realized through a neural network, which approximates the posterior distribution over actions (i.e., ${P}({a}_{t}|{s}_{t})$) using the prior ${P}({a}_{t})$ that is obtained from the MCTS [9]. This network is trained to reproduce the last action sampled form the planner, given the last state. This is similar to the fast and habitual decision-making in biological agents [33].
Fountas et al. (2020) [9] followed the standard four steps of MCTS [34], which let them to restrict and prioritize the trajectories that should be evaluated. Iterativly a weighted tree that has memory updates the visited states. During each loop, a path from the existing tree (towards a leaf node) is selected (i.e., selection) based on the following upper bound confidence:
$U(s_{t},a_{t})=\tilde{G}(s_{t},a_{t})+c_{\mathrm{explore}}\cdot Q_{\phi_{a}}(a_{t}|s_{t})\cdot\frac{1}{1+N(a_{t},s_{t})}$ (6)
@@ 3.3.1 Exploring the Application
The simulation of the system, described in Section 2.1, is designed to replicate features of an industrial system. It utilizes Poisson processes [38] (i.e., exponential distributions) for machine state transitions and part arrivals [7]. The event-driven steps employed by Loffredo et al. (2023) [7, 8] trigger decisions after a system state transition rather than at fixed intervals, proving effective for controlling working machines. This problem can be viewed as either continuous-time stochastic control or a discrete-time Markov Chain process [39]. Continuous-time modeling requires making time intervals visible to the agent for both machines and subsequent observations, whereas discrete-time modeling allows the agent to learn dynamics by observing transitions to create probabilities for different transitions. There is a variable time interval between subsequent observations (i.e., $\Delta t$ as depicted in Fig. 3). This variability requires synchronizing the transition for prediction (i.e., $P_{\theta_{o}}(o_{t+1}|\tilde{s}_{t+1})$) with the next observation (i.e., $\tilde{o}_{t+1}$) in the continuous-time model. Incorporating continuous-time facilitates neural network function approximation for longer $\Delta t$ during planning. However, using continuous-time can further complicate the prediction structure since residence times for machine states exist in observation. Here, we utilize discrete-time event-driven steps that simplify this process compared to the continuous-time approach.
The stochastic nature and the integral and continuous form of the reward functions for the system under study [7, 8] mean decision impacts may not be immediately observable, aligning with POMDPs. The system has a delay in responding to the policy, which we call a long/delayed impact horizon, particularly with respect to reward. This is an important distinction from environments evaluated by Fountas et al. (2020) [9]. The agent proposed by them focuses on planning based on immediate next predictions. The problem at hand is continuous with no terminal state, and the span over which reward functions are integrated may encompass a few thousand steps. This complicates planning and highlights the need for less computational cost.
@@ 3.3.3 Hybrid Horizon
To address the limitations arising from the short horizon of the EFE, which relies on immediate next predictions, we propose augmenting the planner with an auxiliary term to account for longer horizons. Q-learning [41] and its enhanced variant, deep Q-learning, can serve as model-free planners with longer horizons, leveraging rewards to update the Q-value for a state-action pair [42]. Loffredo et al. (2023) [9] demonstrated that deep Q-learning can achieve near-optimal performance for the systems under study. Accordingly, we modify $Q_{\phi_{a}}(a_{t})$ to represent amortized inference of actions, mapping observations $\tilde{o}_{t}$ (or sampled predictions) to normalized action probabilities using a Softmax function and training it with deep Q-learning updates based on rewards coming from experience replay. We introduce a hyperparameter $\gamma$ to balance the contributions of long and short horizons. Thus, we arrive at a new formulation for the planner to incorporate longer horizons:
@@ 4 Results
To demonstrate the potential of our methodology, we concentrate our experiments on controlling a real industrial workstation comprising six parallel-identical machines with an upstream capacity of 10, as outlined in [7]. All the stochastic processes characterizing the workstation follow Poisson distributions, and thus are exponentially distributed with different expected values. We first introduce the preference function, then describe the setup and training, and finally present the numerical performance of our agent [^3] The source code will be available on GitHub at https://github.com/YavarYeganeh/AIF_Meeting_EEC..
@@ 4.1 Preference Function
$R_{\text{production}}$ $=\frac{T_{\text{current}}}{T_{\text{max}}}$ (10a)
$R_{\text{energy}}$ $=1-\frac{E_{\text{avg}}}{E_{\text{max}}}$ (10b)
$R$ $=\phi\cdot R_{\text{production}}+\left(1-\phi\right)\cdot R_{\text{energy}}\ ,$ (10c)
where:
  • $\phi$ is the weighting coefficient balancing production and energy consumption. It is set close to 1 (0.97 in our implementation) to ensure that the agent does not significantly reduce production.
  • $T_{\text{current}}=\frac{NP(t)-NP(t-t_{s})}{t_{s}}$ represents the throughput within the past $t_{s}$ period, where $NP(t)$ is the number of parts produced within this period.
  • $T_{\text{max}}$ is the maximum achievable throughput, occurring under the ALL ON policy [8].
  • $E_{\text{avg}}=\frac{C(t)-C(t-t_{s})}{t_{s}}$ represents the average energy consumption over the past $t_{s}$ period, where $C(t)$ is the total energy consumption within this period.
  • $E_{\text{max}}$ is the maximum theoretical consumption rate of the system, pertaining to all machines operating in their busy state.
@@ 4.2 Agent Setup and Training
We adhere to the MC sampling methodology for calculating EFE, as well as most agent configurations, as outlined by Fountas et al. (2020) [9]. Notably, we employ Bernoulli and Gaussian distributions to model prediction and state distributions, respectively. We introduced a modification by utilizing activation functions for the encoder and transition networks. The output of these two networks generates means and variances representing Gaussian distributions for the state. We applied the Tangent Hyperbolic function for the means and the Sigmoid function, multiplied by a factor, $\lambda_{s}$, between 1 and 2, for the variances. This enforces further regularization and stability for the state space to prevent unbounded values, which need to be fitted into a normal distribution. In our implementation, in contrast to [9], which does not use activations for the state, these activations are essential for learning an effective policy.
@@ 4.3 Agent Performance
This performance is primarily coming from the second term of the EFE (i.e., Eq. 5b), which serves as an intrinsic reward [21], as demonstrated in Fig. 5C. The other two terms are less distinguishable for different actions on average. This suggests that the agent differentiates between various actions in its state space to achieve high rewards. In fact, the long impact horizon and stochasticity, as well as our approximations in transition and planning, hinder the predictive power of the agent, but it managed to infer the impact of different actions in its state space. It is also worth noting that our agent quickly converges to high rewards but may experience instability and loss of control if training continues for (a long time), due to a catastrophic increase in the loss, particularly the reconstruction term of the generative model. This necessitates early stopping mechanisms for training and the introduction of regularization elements (e.g., dropout and normalization) to prevent or mitigate the issue.
@@ 5 Conclusion and Future Work
The results of our study demonstrate the effectiveness of our proposed modifications for the active-inference-inspired agent. Notably, a single multi-transition lookahead with repeated actions, coupled with a hybrid horizon, can achieve high rewards without the need for extensive planning algorithms like MCTS. This is particularly notable considering the unique challenges posed by the application under study, which demands deep lookahead into the future to distinguish different actions, while stochasticity presents a challenge. The significant impact of hyperparameters on agent performance emphasizes the importance of meticulous tuning for optimal results and effective control. This highlights the potential of our methodology to tackle the EEC problem and similar applications characterized by a long/delayed impact horizon and high stochasticity.
Improving the methodology can involve enhancing the generative model, the core of active inference, especially by introducing recurrent transitions or enhancing predictive capabilities. Integration of diffusion-based generative models instead of VAEs [5] is also a promising direction. The framework and formalism of active inference agents show promise for non-stationary scenarios, where model-free agents may struggle to adapt swiftly. Future research will concentrate on improving the agent, extending experimental validation, and tailoring the methodology to non-stationary scenarios, leveraging the strengths of active inference to develop more robust and efficient decision-making algorithms.

### v2

@@ Abstract
Keywords:
Active Inference Probabilistic Deep Learning Reinforcement Learning Energy-Efficient Control Manufacturing Systems
@@ 1 Introduction
Active inference (AIF), an emerging field inspired by the principles of biological brains, offers a promising alternative for decision-making models. It unifies perception, learning, and decision-making under the free energy principle (FEP), which formulates neuronal inference and learning under uncertainty [8]. Accordingly, the brain is modeled through levels of (variational) Bayesian inference [29], minimizing prediction errors by leveraging a generative model of the world while considering uncertainties. This framework enables the development of agents that can calibrate their models and make decisions without complete knowledge of system dynamics. Significant progress has been made in applying active inference across various domains, including robotics, autonomous driving, and healthcare [31, 36, 15], showcasing its ability to handle complex decision-making tasks in dynamic environments. Recently, the manufacturing industry’s focus on energy efficiency has intensified due to its significant contribution to global energy consumption. Addressing energy efficiency at the machine level has become critical, with energy-efficient scheduling (EES) and energy-efficient control (EEC) strategies emerging as key approaches to reducing environmental impact [21]. While traditional EEC methods often necessitate complete system knowledge, reinforcement learning (RL) has shown potential in optimizing manufacturing processes without prior system knowledge [23, 22]. However, RL agents may struggle to rapidly adjust their policies to changing conditions.
@@ 2 Application Overview
2.0.1 EEC in Manufacturing Systems
EEC is gaining prominence in both academic and industrial circles within manufacturing systems. It offers substantial energy savings across three key control levels: component, machine, and production system. At its core, EEC involves managing the power consumption state of objects based on environmental conditions. Objects are kept fully operational when their functions are needed and transitioned to low power states when not in use, though this poses challenges due to the unpredictable nature of demands and the penalties incurred during state transitions. These penalties include both time wasted during transition, when the object adds no value, and the energy consumed during the transition process. A comprehensive and recent literature review on this topic can be found in [33].
2.0.2 System Description
We follow the task outlined by Loffredo et al. (2023) [23], which focuses on a stand-alone manufacturing workstation that can be extended to more complex multi-stage production lines [21]. Accordingly, the system under study has an upstream buffer $B$ with finite holding capacity serving multiple identical parallel machines, as depicted in Fig. 1. It is subject to the stochastic arrival of parts, while machines can transition between various states: working (idle or busy), standby, startup, and failed. Power consumption varies by state: standby ($w_{sb}$), failed ($w_{f}$), startup ($w_{su}$), idle ($w_{id}$), and busy ($w_{b}$), where $w_{b}>w_{su}>w_{id}>w_{sb}\approx w_{f}\approx 0$.
Figure 1: Layout of parallel and identical machines in the workstation [23].
The key characteristic of this system is that all processes are stochastic, modeled as Poisson processes [19]. This pertains to the arrival rate ($\lambda$) to buffer $B$, machine processing times ($\mu$), startup times ($\delta$), time between failures ($\psi$), and time to repair ($\xi$), all with expected values, independent and stationary. Machines work on a single part type following a first-come-first-served rule and cannot be switched off during processing or startup. Machines become starved (i.e., enter an idle state) if they are ready to process but $B$ is empty, and they cannot be blocked as there is an infinite downstream buffer. Effective EEC aims to manage power variations efficiently to minimize energy consumption without compromising productivity. Therefore, the primary task is to dynamically decide the number of machines to keep working or switch off based on the emergent stochastic patterns, ensuring optimal trade-offs between energy savings and system throughput.
@@ 3.1 Active Inference Formalism
where $\sigma(\cdot)$ represents the Softmax function, and $\pi$ (i.e., policy) denotes the sequence of actions. The EFE encompasses minimizing surprise regarding preferred observations  [^1] In EFE, the surprise of predictions is measured with respect to the preference, while in VFE, it is measured with respect to the actual observation used to calibrate the model.,exploring uncertainty, and reducing uncertainty about model parameters [7]. The EFE for $\tau\geq t$ can be formulated as follows [37]:
@@ 3.2 Architecture
An agent within the active inference framework needs different modules, which are entangled within the framework. Amortization [18, 25, 12] is introduced into the formalism (in Sec. 3.1) to scale-up the realization [7]. The formalism is inherently probabilistic and parameterized with with two sets, $\theta=\{{\theta}_{s},{\theta}_{o}\}$ representing generative and $\phi=\{{\phi}_{s}\}$ recognition elements [7]. Using the following parameterized modules the formalism (in Sec. 3.1) can be calculated: Encoder (i.e., $Q_{\phi_{s}}(s_{t})$), an amortized inference of the hidden state (i.e., an inference network [24] providing a mapping between the observation, $\tilde{o}_{t}$, and a distribution for its corresponding hidden state). Transition (i.e., ${P}_{\theta_{s}}({s}_{t+1}|\tilde{s}_{t},\tilde{a}_{t})$), which generates a distribution for the next hidden state based on both a sampled action and the current hidden state. Decoder (i.e., ${P}_{\theta_{o}}({o}_{t+1}|\tilde{s}_{t+1})$) generates a distribution for prediction based on the sampled hidden state.
@@
Utilizing the mentioned architecture (also depicted in Fig. 3), the agent would be able to calculate the EFE in Eq. 5 for a given policy (i.e., $\pi$), which is a sequence of actions. Thus, we can calculate the probabilities for selecting actions through Eq. 2. In order to make better decisions, the agent can plan ahead by simulating future trajectories using the architecture mentioned above. However, as the policy space grows exponentially into the future, it is infeasible to evaluate all possible scenarios. Fountas et al. (2020) [7] introduced two approaches to alleviate the obstacle. 1) They used the standard Monte-Carlo Tree Search (MCTS) [3, 38], a tree-based search algorithm, which selectively explore promising trajectories in a restricted manner. 2) They introduced another recognition module [32, 25, 41], parameterized with ${\phi}_{a}$ as follows: Habit (i.e., $Q_{\phi_{a}}(a_{t})$), an amortized inference of actions (i.e., an inference network [24] providing a mapping between a sampled hidden state, $\tilde{s}_{t}$, and normalized, through Softmax function, probabilities for the actions).
@@
Habit is also realized through a neural network, which approximates the posterior distribution over actions (i.e., ${P}({a}_{t}|{s}_{t})$) using the prior ${P}({a}_{t})$ that is obtained from the MCTS [7]. This network is trained to reproduce the last action sampled form the planner, given the last state. This is similar to the fast and habitual decision-making in biological agents [42].
Fountas et al. (2020) [7] followed the standard four steps of MCTS [38, 40], which let them to restrict and prioritize the trajectories that should be evaluated. Iterativly a weighted tree that has memory updates the visited states. During each loop, a path from the existing tree (towards a leaf node) is selected (i.e., selection) based on the following upper bound confidence:
$U(s_{t},a_{t})=\tilde{G}(s_{t},a_{t})+c_{\mathrm{explore}}\cdot Q_{\phi_{a}}(a_{t}|s_{t})\cdot\frac{1}{1+N(a_{t},s_{t})}\ .$ (6)
@@ 3.3.1 Exploring the Application
The simulation of the system described in Section 2.0.1, is designed to replicate features of an industrial system. It utilizes Poisson processes [19] (i.e., exponential distributions) for machine state transitions and part arrivals [23]. The event-driven steps employed by Loffredo et al. (2023) [23, 22] trigger decisions after a system state transition rather than at fixed intervals, proving effective for controlling working machines. This problem can be viewed as either continuous-time stochastic control or a discrete-time Markov Chain process [34]. Continuous-time modeling requires making time intervals visible to the agent for both machines and subsequent observations, whereas discrete-time modeling allows the agent to learn dynamics by observing transitions to create probabilities for different transitions. There is a variable time interval between subsequent observations (i.e., $\Delta t$ as depicted in Fig. 3). This variability requires synchronizing the transition for prediction (i.e., $P_{\theta_{o}}(o_{t+1}|\tilde{s}_{t+1})$) with the next observation (i.e., $\tilde{o}_{t+1}$) in the continuous-time model. Incorporating continuous-time facilitates neural network function approximation for longer $\Delta t$ during planning. However, using continuous-time can further complicate the prediction structure since residence times for machine states exist in observation. Here, we utilize discrete-time event-driven steps that simplify this process compared to the continuous-time approach.
The stochastic nature, along with the integral and continuous form of the reward functions for the system under study [23, 22], implies that the effects of decisions may not be immediately observable, aligning with POMDPs. The system has a delay in responding to the policy (i.e., delayed policy response), which we refer to as long/delayed impact horizon, particularly with respect to reward. This is an important distinction from environments evaluated by Fountas et al. (2020) [7]. The agent proposed by them focuses on planning based on immediate next predictions. The problem at hand is continuous with no terminal state, and the span over which reward functions are integrated may encompass a few thousand steps. This complicates planning and highlights the need for less computational cost.
@@ 3.3.3 Hybrid Horizon
To address the limitations arising from the short horizon of the EFE, which relies on immediate next predictions, we propose augmenting the planner with an auxiliary term to account for longer horizons. Q-learning [43] and its enhanced variant, deep Q-learning [28], can serve as model-free planners with longer horizons, leveraging rewards to update the Q-value for a state-action pair. The Q-value represents the expected return, considering long-term consequences even in one-step lookahead [39]. Mnih et al. (2015) [28] demonstrated the effectiveness of DQN in learning relatively long-term strategies in certain games. Loffredo et al. (2023) [23] demonstrated that deep Q-learning can achieve near-optimal performance for the systems under study. Accordingly, we modify $Q_{\phi_{a}}(a_{t})$ to represent amortized inference of actions, mapping observations $\tilde{o}_{t}$ (or sampled predictions) to normalized action probabilities using a Softmax function and training it with deep Q-learning updates based on rewards from experience replay. We introduce a hyperparameter $\gamma$ to balance the contributions of long and short horizons. Thus, we arrive at a new formulation for the planner to incorporate longer horizons:
@@ 4 Results
To demonstrate the potential of our methodology, we concentrate our experiments on controlling a real industrial workstation comprising six parallel-identical machines with an upstream capacity of 10, as outlined in [23]. All the stochastic processes characterizing the workstation follow Poisson distributions, and thus are exponentially distributed with different expected values. We first introduce the preference function, then describe the setup and training, and finally present the numerical performance of our agent [^3] The source code is available at https://github.com/YavarYeganeh/AIF_Meeting_EEC..
@@ 4.1 Preference Function
$R_{\text{production}}$ $=\frac{T_{\text{current}}}{T_{\text{max}}}\quad R_{\text{energy}}=1-\frac{E_{\text{avg}}}{E_{\text{max}}}$ (10a)
$R$ $=\phi\cdot R_{\text{production}}+\left(1-\phi\right)\cdot R_{\text{energy}}\ ,$ (10b)
where $\phi$ is the weighting coefficient balancing production and energy consumption. $\phi$ is set close to 1 (0.97 in our implementation) to ensure that the agent does not significantly reduce production. $T_{\text{current}}=\frac{NP(t)-NP(t-t_{s})}{t_{s}}$ represents the throughput within the past $t_{s}$ period, where $NP(t)$ is the number of parts produced within this period. $T_{\text{max}}$ is the maximum achievable throughput, occurring under the ALL ON policy [22]. $E_{\text{avg}}=\frac{C(t)-C(t-t_{s})}{t_{s}}$ represents the average energy consumption over the past $t_{s}$ period, where $C(t)$ is the total energy consumption within this period. $E_{\text{max}}$ is the maximum theoretical consumption rate of the system, pertaining to all machines operating in their busy state.
@@ 4.2 Agent Setup and Training
We adhere to the MC sampling methodology for calculating EFE, as well as most agent configurations, as outlined by Fountas et al. (2020) [7]. Notably, we employ Bernoulli and Gaussian distributions to model prediction and state distributions, respectively. We introduced a modification by utilizing activation functions for the encoder and transition networks. The output of these two networks generates means and variances representing Gaussian distributions for the state. We applied the Tangent Hyperbolic function for the means and the Sigmoid function, multiplied by a factor, $\lambda_{s}$, between 1 and 2, for the variances. This enforces further regularization and stability for the state space to prevent unbounded values, which need to be fitted into a normal distribution. In contrast to [7], which does not use activations for the state, our implementation includes these activations, as we found them essential for learning effective policies to achieve high rewards.
@@ 4.3 Agent Performance
This performance is primarily coming from the second term of the EFE (i.e., Eq. 5b), which serves as an intrinsic reward [30], as demonstrated in Fig. 5C. The other two terms are less distinguishable for different actions on average. This suggests that the agent differentiates between various actions in its state space to achieve high rewards. In fact, the long impact horizon and stochasticity, as well as our approximations in transition and planning, hinder the predictive power of the agent, but it managed to infer the impact of different actions in its state space. It is also worth noting that our agent quickly converges to high rewards but may experience instability and loss of control if training continues for (a long time), due to a catastrophic increase in the loss, particularly the reconstruction term of the generative model. This necessitates early stopping mechanisms for training and the introduction of regularization elements [13] (e.g., dropout and normalization) to prevent or mitigate the issue.
@@ 5 Conclusion and Future Work
The results demonstrate the effectiveness of our proposed modifications for the active-inference-inspired agent. Notably, a single multi-transition lookahead with repeated actions, coupled with a hybrid horizon, achieves high rewards without relying on extensive planning algorithms like MCTS. This is important given the application’s need for deep lookaheads into the future, while stochasticity presents a challenge. Unlike RL, which generally depends on high-quality reward signals that can be expensive to collect and difficult to design, AIF operates directly on observations. This reduces the need for well-engineered and expensive reward structures while enabling more flexible and adaptive decision-making. Overall, the potential of our methodology for addressing the EEC problem and similar scenarios characterized by delayed policy response and high stochasticity is evident.
Improving the methodology could involve enhancing the generative model, the core of active inference, especially by introducing recurrent transitions or enhancing predictive capabilities. The integration of diffusion-based generative models instead of VAEs [15] is also a promising direction. The framework and formalism of active inference agents show promise for non-stationary scenarios, where model-free agents may struggle to adapt swiftly. Future research will focus on improving the agent, extending experimental validation, and tailoring the methodology for non-stationary scenarios, leveraging the strengths of active inference to develop more robust and efficient decision-making algorithms.
Acknowledgments
The work presented in this paper was supported by HiCONNECTS, which has received funding from the Key Digital Technologies Joint Undertaking under grant agreement No. 101097296.